(a) List all the generators of < 5 > in Z6o. (b) List all of the left cosets of < 10 > in the subgroup < 2 > of Z60.
Answer #1

Date: 11 (25) = 0 of H. = 45 a = Zoo, HSG e, H= <s> M-Zos,10,15,20, 25, 30, 35,40,45,50, 55} Now 0(4)=12 generators of t have2 0 (45) 0(x)= 12 12 = 4 붕 gd(9,12) 050)Ol10x5) 12 12-6 ged(10, 12) 0(56)= olix)= 123 12 = 12 V ged(11,12) 1 : elements of ornow 20 +K = K x 26eK 2otK= К А Я ОК 30 + К= КщЗоєК ЧО+KE K as 40EK 50+ K = as soek :. - 4K, 2-к, 4+K, +6 +k, 8+к, 12+к, 14+K,